The lifetime of a certain kind of bulb has a normal distribution with mean of 500 hours and standard deviation of 45 hours

Find the percentage of bulbs with a lifetime of at least 570 hrs

The percentage of bulbs with lifetime in between 485 and 515 hrs

The minimum lifetime of the best 5% of the bulbs


1
Expert's answer
2022-02-21T15:02:33-0500

Let "X=" lifetime of a certain kind of bulb: "X\sim N(\mu, \sigma^2)."

Given "\mu=500h, \sigma=45h."

i)


"P(X\ge570)=1-P(X<570)"

"=1-P(Z<\dfrac{570-500}{45})"

"\approx1-P(Z<1.555556)\approx0.0599"

"5.99\%"


ii)



"P(485<X<515)=P(X<515)-P(X\le485)"

"=P(Z<\dfrac{515-500}{45})-P(Z\le\dfrac{485-500}{45})"

"\approx P(Z<0.333333)-P(Z<-0.333333)\approx0.2611"

"26.11\%"


iii)


"P(X\ge x)=1-P(X<x)"

"=1-P(Z<\dfrac{x-500}{45})=0.05"

"\dfrac{x-500}{45}\approx1.6449""x=45(1.6449)+500"

"x=574\ hours"